# Memorization-Compression Cycles Improve Generalization

### v1

@@ 1 Introduction
Note that $H(Y|\hat{Y})$ is the cross-entropy (CE) loss in conventional language modeling task [11]. In short, LBLM requires maximal compressing internal representation under optimal cross-entropy. To explicitly calculate $H(R)$, we adopt Matrix-based entropy (MBE), first proposed in [5], given a matrix $R\in\mathbb{R}^{s\times d}$ concatenating $s$ token representations, MBE is given by
@@
where $K=RR^{T}$ is the Gram matrix. MBE essentially provide a continuous measure of matrix rank, by calculating entropy on distribution of singular values. It has been shown to exhibit a strong correlation with embedding quality [14], where it’s also observed that later checkpoints in pre-trained base models has lower entropy, suggesting pre-training with cross-entropy loss alone leads to decrease in MBE.
Previous work [15] observed that in deterministic models (where $I(R;X)=H(R)$), training with a single loss target leads to a two-phase trend: a short initial memorization phase with rapid empirical loss reduction, followed by a monotonic decrease in $H(R)$, interpreted as a compression phase. This observation suggests a clean separation between learning and compression in deep networks.
However, our empirical analysis of GPT pretraining reveals a richer structure. By tracking the cosine similarity between gradients of CE and MBE, we observe that their alignment oscillates between positive and negative throughout training. This suggests that instead of proceeding through distinct phases, learning unfolds as a cyclic process, continuously alternating between expanding representations to capture data (memorization) and reorganizing them into more compact forms (compression). Importantly, this local oscillation occurs alongside a global trend of decreasing MBE, indicating that compression accumulates over time even as the network periodically re-enters memorization-like states. We refer to this phenomenon as the memorization–compression cycle.
In biological neural systems, learning is inherently cyclic, alternating between awake learning and sleep consolidation. In a seminal study [18], sleep was shown to resolve interference between conflicting memories by encouraging competition between them. When two conflicting tasks were presented sequentially, awake learning alone led to catastrophic forgetting, as new synaptic updates overwrote previously learned associations. Sleep consolidation overcame this by reorganizing synaptic weights—allocating distinct subsets to store each memory—achieving better retention than even interleaved learning.
Inspired by biolocial learning cycle, we introduce Gated Phase Transition (GAPT), a training algorithm that explicitly alternates between memorization (minimizing CE) and compression (minimizing CE and MBE) phases. GAPT is designed to approximate the optimal solution to the IBLM objective by dynamically switching phases based on learning signals, mirroring the role of consolidation in resolving representational conflict.
GAPT delivers consistent improvements across domains:
First, in LLM pre-training on FineWeb dataset, GAPT reduces Matrix-Based Entropy (MBE) by an average of $70.5\%$ across layers while improving cross-entropy by $4.8\%$, outperforming standard language modeling and approaching the IBLM trade-off.
Second, in arithmetic generalization, GAPT reduces test cross-entropy by $35\%$ and MBE by $47\%$ when trained on 1–3 digit multiplication and tested on 4–6 digits, supporting Theorem 1.
Third, in a synthetic task with gradient interference, GAPT improves representational separation by 97% and reduces MBE by $91\%$ relative to a mixed baseline—closely mirroring the conflict resolution behavior observed in biological sleep.
@@
The remainder of the paper expands on these contributions: Section 2 presents our theoretical framework and generalization bound; Section 3 details the GAPT algorithm; Section 4 presents empirical results, including the memorization–compression cycle and GAPT’s effectiveness; Section 5 discusses relevant work; Section 6 discusses broader implications.
@@ 2 Theory
where $|\mathcal{X}|$ denotes the cardinality of the input space. By Corollory 1, we have $|\mathcal{X}|=\mathcal{O}(2^{\alpha\cdot H(X)})$ for $\alpha\in[1,+\infty)$ , this suggests the upper bound could be reformulated via entropy
$\mathcal{L}_{P(X,Y)}(f,\ell)-\hat{\mathcal{L}}_{P(X,Y)}(f,\ell,\mathcal{D}_{N})\leq\mathcal{O}(\frac{2^{\alpha\cdot H(X)}\log N}{\sqrt{N}})$
Let $f$ be a neural network with $L$ intermediate representations forming a Markov chain:
@@
Fix any intermediate representation $R_{l}$, and decompose the network into $f=d\circ e$, where $e:\mathcal{X}\rightarrow\mathcal{R}_{l}$ maps inputs to $R_{l}=e(X)$, and $d:\mathcal{R}_{l}\rightarrow\mathcal{Y}$ predicts the output.
Applying the generalization bound to the representation $R_{l}$, we obtain:
@@
^$\leq\hat{\mathcal{L}}_{P(R_{l},Y)}(d,\ell_{e},\mathcal{D}_{N})+\mathcal{O}\left(\frac{2^{\alpha\cdot H(R_{l})}\log N}{\sqrt{N}}\right)$ (4)
^$\leq\hat{\mathcal{L}}_{P(X,Y)}(f,\ell,\mathcal{D}_{N})+\mathcal{O}\left(\frac{2^{\alpha\cdot H(R_{l})}\log N}{\sqrt{N}}\right)$ (5)
@@
The Information Bottleneck (IB) framework [8] defines the optimal representation $R$ as the solution to:
@@
subject to $I(Y;R)=I(Y;X)$ (7)
Intuitively, the optimal $R$ is one that discards as much information from the input $X$ as possible, while retaining all information necessary to predict the output $Y$. This representation is referred to as the minimal sufficient statistic.
In the case of large language models (LLMs), the network forms a deterministic Markov chain:
$X\rightarrow R\rightarrow\hat{Y}$
where $R$ denotes an intermediate hidden representation and $\hat{Y}$ is the predicted output.
Because $R$ is deterministically computed from $X$, we have:
$I(X;R)$ $=H(R)-H(R|X)=H(R)$ (8)
$I(Y;X)$ $\geq I(Y;R)\geq I(Y;\hat{Y})$ (9)
$I(Y;\hat{Y})$ $=H(Y)-H(\hat{Y}|Y)$ (10)
The first equality follows from the fact that $H(R|X)=0$ for deterministic functions. The second line follows from the Data Processing Inequality (DPI) [3], which ensures that information cannot increase along a Markov chain.
Rewriting the loss of predictive information:
$I(Y;X)-I(Y;R)$ $\leq I(Y;X)-I(Y;\hat{Y})$ (11)
$=H(Y|\hat{Y})-H(Y|X)$ (12)
Since $H(Y|X)$ is fixed (defined by the true distribution), minimizing $H(Y|\hat{Y})$ increases $I(Y;\hat{Y})$ and pushes it closer to $I(Y;X)$. In practice, minimizing $H(Y|\hat{Y})$ corresponds to minimizing the cross-entropy loss used in language modeling.
On the compression side, we know $H(R|X)=0$ since neural network propagation is deterministic, therefore we have $I(X;R)=H(R)$. Thus, minimizing $H(R)$ directly minimizes the IB compression term $I(X;R)$, giving us a tractable surrogate for representation compression.
Together, this shows that minimizing cross-entropy corresponds to satisfying the predictive constraint in IB, while minimizing representation entropy implements the compression objective—justifying the IBLM formulation. ∎
@@ 4.1 Natural Compression-Memorization Cycle
In Figure 1, we observe that training with CE loss alone leads to a consistent decrease in MBE across layers, confirming observations from [14]. However, unlike the "compression valley" phenomenon described in that work, our results show that layers 7–12 have lower MBE than layers 1–6. This difference is likely due to our architectural design: we adopt a U-Net-like skip connection pattern in which layer 1 is connected to layer 12, layer 2 to layer 11, and so on. As a result, the lower MBE observed in later layers likely reflects more compact representations in the decoding stages.
@@
We further analyze CE gradient behavior. Figure 2. shows that gradient consistency (measured by cosine similarity across batches) declines over time across all layers. This indicates an increasing signal-to-noise ratio in CE gradients as training progresses.
@@
We next inspect cosine similarity between CE and MBE gradients. As shown in Figure 3, we observe recurring sign flips that indicate an implicit alternation between memorization (negative similarity) and compression (positive similarity) phases, even without explicitly optimizing for entropy.
To characterize this oscillation, we analyze the gradient signal across training using three measures: (1) standard deviation (oscillation strength), (2) zero-crossing rate (oscillation frequency), and (3) peak-to-average power ratio from the power spectral density (periodicity strength). Figure 4 summarizes these metrics across layers and parameter types. We find that attention parameters exhibit stronger and more frequent oscillations than MLP parameters. Interestingly, earlier layers show higher oscillation frequency, while no layer demonstrates strong rhythmic periodicity—suggesting that oscillation is irregular and state-driven, rather than strictly periodic.
This suggests an intriguing perspective on learning: while training on a fixed dataset induces a global momentum toward representation compression, the local dynamics alternate between phases of memorization and compression.
@@ 4.2 Pre-training with GAPT
In our second experiment, we retain the same GPT pre-training setup but incorporate the GAPT algorithm to test whether it offers a better solution to IBLM objective defined in Equation 2 , compared to a baseline model trained solely on the cross-entropy (CE) loss. MBE regularization during compression phase in GAPT is done from layer 2 to 9.
@@
Table 1: Cross-entropy loss on the FineWeb validation set.
As shown in Table 1, GAPT reduces cross-entropy on the FineWeb validation set by $4.8\%$ compared to the CE-only baseline. In addition to reducing test CE loss, GAPT also significantly compresses internal representations. Figure 5 (left) shows the layer-wise MBE for both models. We observe consistent reductions across all regularized layers. The per-layer MBE values and their relative improvements are summarized in Figure 5 (right).
GAPT reduces MBE by an average of $70.5\%$ across layers 2–9 while improving validation cross-entropy by $4.8\%$. This suggests that explicitly alternating between memorization and compression phases offers an effective solution to the constrained optimization objective in IBLM (Equation 2), and can exceed baseline training with cross-entropy target alone.
@@ 4.4 Arithmetic Generalization
We tokenize the input and output sequences at the per-digit level and train the model for 1,750 iterations with a batch size of 16. Due to observed instability in OOD entropy, we adopt an early stopping strategy that halts training if validation loss increases by more than $20\%$ after iteration 800. We expect future work to further stabilize GAPT in OOD settings.
@@ 5 Relevant Work
The Information Bottleneck (IB) method was introduced in [2] to formalize the goal of retaining only task-relevant information in representations by minimizing entropy. A theoretical connection between generalization and the cardinality of discrete inputs was established in [6], and later applied to deep networks in [8]. However, the entropy–generalization link remained incomplete due to the gap between cardinality and entropy measures.
Empirical evidence of a two-phase learning dynamic—early memorization followed by entropy compression—was presented in [15]. To quantify entropy in neural networks, matrix-based entropy (MBE) was proposed in [5], and later applied to LLMs in [14], where it correlated with embedding quality and revealed compression trends across checkpoints.
LLMs such as GPT-2 [11] generalize across tasks by scaling training data and parameters [4], effectively compressing the training corpus [17, 13]. While post-training methods like instruction tuning [9] improve usability, LLMs still struggle on out-of-distribution tasks such as reverse reasoning [22] and multi-hop inference [20].
Recent advancements in RL with verifiable rewards (RLVR) have improved mathematical and coding performance [7], though often by narrowing base model behavior [10]. Finally, vocabulary design has seen progress through curriculum-based tokenization, where dynamic vocabulary growth based on modeling entropy leads to improved pretraining efficiency [21].
6 Implication
Generalization occurs when learning from one task improves performance on another. Our findings suggest that generalization can emerge from compressing internal representations under predictive constraints—without relying on massive amounts of demonstrative data. This ability may allow artificial systems to discover novel patterns and principles, reducing dependence on large-scale supervision. However, it also raises important safety concerns: a system that generalizes well but lacks transparency or interpretability may pose greater risks than one that mostly memorizes. As we move toward AGI, it is essential to understand and govern the mechanisms behind representation compression and generalization to ensure safe, aligned, and accountable behavior.

### v2

# Non-Euclidean Foundation Models and Geometric Learning

@@ 1 Introduction
[…]
@@
where $K=RR^{T}$ is the Gram matrix. MBE essentially provide a continuous measure of matrix rank, by calculating entropy on distribution of singular values. It has been shown to exhibit a strong correlation with embedding quality [14].
Previous work [15] showed that in neural networks, training often follows a two-phase trend: an initial memorization phase with rapid loss reduction, followed by a monotonic decrease in representation entropy $H(R)$, interpreted as compression. This suggests a clean, sequential separation between fitting and compressing.
Our analysis of GPT pretraining reveals a richer dynamic: the cosine similarity between CE and MBE gradients oscillates between positive and negative, indicating a cyclic alternation between memorization (representation expansion) and compression. Rather than progressing through distinct phases, learning continuously revisits these states. Notably, this local oscillation coexists with a global decline in MBE, suggesting that compression accumulates over time even as the model periodically re-engages in memorization. We term this the memorization–compression cycle.
In biological neural systems, learning proceeds cyclically, alternating between active acquisition and sleep-driven consolidation. A seminal study [18] showed that sleep mitigates interference between conflicting memories by compressing and reorganizing them—allocating distinct subsets of synaptic weights to each. When two conflicting tasks were presented sequentially, awake learning alone led to catastrophic forgetting, while sleep consolidation enabled the formation of separable representations, outperforming even interleaved training. Inspired by this, we propose Gated Phase Transition (GAPT), a training algorithm that alternates between memorization (minimizing cross-entropy) and compression (minimizing cross-entropy and Matrix-Based Entropy). GAPT dynamically switches phases based on learning signals, approximating the IBLM objective and functionally mirroring biological consolidation to resolve representational conflict.
[…]
@@
The remainder of the paper expands on these contributions: Section 2 presents our theoretical framework and generalization bound; Section 3 details the GAPT algorithm; Section 4 presents empirical results, including the memorization–compression cycle and GAPT’s effectiveness; Section 5 discusses relevant work.
@@ 2 Theory
$\leq\mathcal{O}(\frac{2^{\alpha\cdot H(X)}\log N}{\sqrt{N}})$
where $|\mathcal{X}|$ denotes the cardinality of the input space. Second line follows from Corollory 1, which states that we have $|\mathcal{X}|=\mathcal{O}(2^{\alpha\cdot H(X)})$ for $\alpha\in[1,+\infty)$. Let $f$ be a neural network with $L$ intermediate representations forming a Markov chain:
@@
Fix any intermediate representation $R_{l}$, and decompose the network into $f=d\circ e$, where $e:\mathcal{X}\rightarrow\mathcal{R}_{l}$ maps inputs to $R_{l}=e(X)$, and $d:\mathcal{R}_{l}\rightarrow\mathcal{Y}$ predicts the output. Applying the generalization bound to the representation $R_{l}$, we obtain:
@@
$\leq\hat{\mathcal{L}}_{P(R_{l},Y)}(d,\ell_{e},\mathcal{D}_{N})+\mathcal{O}\left(\frac{2^{\alpha\cdot H(R_{l})}\log N}{\sqrt{N}}\right)$ (4)
$\leq\hat{\mathcal{L}}_{P(X,Y)}(f,\ell,\mathcal{D}_{N})+\mathcal{O}\left(\frac{2^{\alpha\cdot H(R_{l})}\log N}{\sqrt{N}}\right)$ (5)
@@
The Information Bottleneck (IB) framework [2, 8] defines the optimal representation $R$ as the solution to:
@@
s.t. $I(Y;R)=I(Y;X)$ (7)
That is, $R$ should discard as much input information as possible while preserving all information relevant to predicting YY—the minimal sufficient statistic. In large language models (LLMs), the network forms a deterministic Markov chain: $X\rightarrow R\rightarrow\hat{Y}$. Since $R$ is deterministically computed from $X$, we have:
$I(X;R)$ $=H(R)\quad\text{(as }H(R|X)=0\text{)}$ (8)
[…]
Where second line follows from data processing inequality [3]. Thus,
$I(Y;X)-I(Y;R)$ $\leq H(Y|\hat{Y})-H(Y|X)$ (10)
[…]
@@ 4.1 Natural Compression-Memorization cycle
In Figure 1, we observe that training with CE loss alone leads to a consistent decrease in MBE across layers, confirming observations from [14]. We further analyze CE gradient behavior. Figure 2. shows that gradient consistency (measured by cosine similarity across batches) declines over time across all layers. This indicates an increasing signal-to-noise ratio in CE gradients as training progresses.
@@
We next inspect cosine similarity between CE and MBE gradients. As shown in Figure 3, we observe recurring sign flips that indicate an alternation between memorization and compression phases, even without explicitly optimizing for entropy. To characterize this oscillation, we analyze the gradient signal across training using three measures: standard deviation (oscillation strength), zero-crossing rate (oscillation frequency), and peak-to-average power ratio from the power spectral density (periodicity strength). Figure 4 summarizes these metrics across layers and parameters. We find that attention parameters exhibit stronger and more frequent oscillations than MLP parameters. Interestingly, later layers show higher oscillation frequency, while no layer demonstrates strong rhythmic periodicity—suggesting that oscillation is irregular and state-driven, rather than strictly periodic.
@@ 4.2 Pre-training with GAPT
In our second experiment, we retain the same GPT pre-training setup but incorporate the GAPT algorithm to test whether it offers a better solution to IBLM objective defined in Equation 2 , compared to a baseline model trained solely on the CE loss. MBE regularization during compression phase in GAPT is done from layer 2 to 9.
Table 1: Cross-entropy loss on the FineWeb validation set.
@@
As shown in Table 1, GAPT reduces cross-entropy on the FineWeb validation set by $4.8\%$ compared to the CE-only baseline. In addition to reducing test CE loss, GAPT also significantly compresses internal representations. Figure 5 (left) shows the layer-wise MBE for both models. We observe consistent reductions across all regularized layers. The per-layer MBE values and their relative improvements are summarized in Figure 5 (right). GAPT reduces MBE by an average of $70.5\%$ across layers 2–9 while improving validation cross-entropy by $4.8\%$. This suggests that explicitly alternating between memorization and compression phases offers an effective solution to the constrained optimization objective in IBLM (Equation 2), and can exceed baseline training with cross-entropy target alone.
@@ 4.4 Arithmetic generalization
We tokenize the input and output sequences at the per-digit level and train the model for 1,750 iterations with a batch size of 16. Due to observed instability in OOD entropy, we adopt an early stopping strategy that halts training if validation loss increases by more than $20\%$ after iteration 800.
@@ 5 Relevant work
The Information Bottleneck (IB) method was introduced in [2] to formalize the goal of retaining only task-relevant information in representations by minimizing entropy. A theoretical connection between generalization and the cardinality of discrete inputs was established in [6], and later applied to deep networks in [8]. However, the entropy–generalization link remained incomplete due to the gap between cardinality and entropy measures. Empirical evidence of a two-phase learning dynamic—early memorization followed by entropy compression—was presented in [15]. To quantify entropy in neural networks, matrix-based entropy (MBE) was proposed in [5], and later applied to LLMs in [14], where it correlated with embedding quality and revealed compression trends across checkpoints.
LLMs such as GPT-2 [11] generalize across tasks by scaling training data and parameters [4], effectively compressing the training corpus [17, 13]. While post-training methods like instruction tuning [9] improve usability, LLMs still struggle on out-of-distribution tasks such as reverse reasoning [22, 20] and multi-hop inference [20]. Recent advancements in RL with verifiable rewards (RLVR) have improved mathematical and coding performance [7], though often by narrowing base model behavior [10]. Finally, curriculum-based tokenization growth has been combined with pre-training to improve compression [21].
@@ Appendix A Appendix
Broader Impact and Limitations
Limitations.
While theoretically grounded, our proposed GAPT algorithm remains relatively simple. On large-scale pretraining tasks, it achieves less than a 5% improvement over the baseline, falling short of fully realizing the theoretical benefits of compact representations for generalization as suggested by Theorem 1. Furthermore, scaling experiments are missing, which requires more computation resources but may realize more potential in this new dimension for improving generalization. Additionally, entropy minimization is only one approach to compression; other mechanisms—such as sparse activation—may offer stronger biological plausibility and efficiency. A further limitation is observed in the arithmetic generalization task, where GAPT exhibits instability in out-of-distribution performance across different training runs. There could be some data selection side piece missing for stabilizing generalization-oriented training runs.
Broader Impact.
Enhancing generalization in AI systems holds the potential to enable models that can extrapolate, reason, and discover novel knowledge—key milestones toward artificial general intelligence. However, systems that generalize without explainability may pose greater risks than those that merely memorize. We caution that aggressive compression of internal representations, while beneficial for efficiency and abstraction, may also increase susceptibility to unintended behaviors such as spurious generalization or adversarial misuse.